OpenStudy (anonymous):

what is the quotient of 3/43 divided by 25/86

OpenStudy (anonymous):

\[\large \frac{3}{43}\times \frac{86}{25}\] is a start cancel first, multiply last

OpenStudy (blackops2luvr):

@kaitlin_nikole Don't jut give the answers out

OpenStudy (anonymous):

sorry.